Question 1 Construct a 95% confidence interval for the average value of y for the following data. x 12 21 288 20 y 17 15 22 19 24 Use x = 25, se 3.94, and the equation of the regression line, y = 16.5 + 0.162x Do not round the intermediate values. Round your answers to 2 decimal places.)
Answer #1

Ans:

degree of freedom=5-2=3

critical t value=tinv(0.05,3)=3.182

y'(25)=16.5+0.162*25=20.55

95% confidence interval for average value of y

=20.55+/-3.182*3.94

=20.55+/-12.54

=(8.01,33.09)